OpenStudy (anonymous):

PLEASE HELP Consider the three points P1 : (1; 1; 0), P2 : (6; 2; 1), and P3 : (4; 3; 2). (a) [1 marks] Find the displacement vector from P1 to P2. (b) [1 marks] Find the displacement vector from P1 to P3. (c) [4 marks] Find the angle between the two vectors found in parts (a) and (b).

OpenStudy (anonymous):

1. P1 - P2 2. P1 - P3 \[3. \cos \theta = \frac{ ||a||\,\,||b|| }{ a\cdot b }\text{ where }a\cdot b \text{ is the inner product}\]

OpenStudy (anonymous):

so with p1-p2 would it be lthe answer be(1-6, 1-2, 0-1)?

OpenStudy (anonymous):

P1-P2 = (-5, -1, -1)

OpenStudy (anonymous):

wait but isnt rule that you go p2-pi because you go final minus initial?

OpenStudy (anonymous):

"displacement vector from P1 to P2"

OpenStudy (anonymous):

yeah so p2 is the final point and p1 is the initial point. P1 to P2

OpenStudy (anonymous):

yeah it would be... to go from 1 to 6 you have to add 5, etc...

OpenStudy (anonymous):

oh sorry I had typed the question in wrong. it was -6 in P2 and -2 in P3. My bad

OpenStudy (anonymous):

it okay... i flipped the formula for cos...\[\cos \theta =\frac{ a\cdot b }{ ||a||\,\,||b|| }\]

OpenStudy (anonymous):

the formula you had originally wsa right because a.b = |a||||b|| cosθ

OpenStudy (anonymous):

yes but solve for cos

OpenStudy (anonymous):

man i am out of it today

OpenStudy (anonymous):

no... the last is correct not the first

OpenStudy (anonymous):

Stupid mistake thanks for your help

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

I get it

OpenStudy (anonymous):

sleep

OpenStudy (anonymous):

ha will do

OpenStudy (anonymous):

for me!

OpenStudy (anonymous):

woops well sleep for the both of us then

OpenStudy (anonymous):

lol